OpenStudy (anonymous):

The perimeter of a rectangular yardis 270 feet. if its length is 25 feet greater than its width, what are the dimensions of the yard? help!

OpenStudy (kropot72):

Let width = w Then Length = w + 25 Perimeter = (2 * w) + (2 * (w + 25)= 270 4w + 50 = 270 Subtract 50 from each side: 4w = 220 w = 55 feet Dimensions of yard are: 55 feet by 80 feet

OpenStudy (anonymous):

thank you sooo much!!!

OpenStudy (kropot72):

You're welcome :)